OpenStudy (anonymous):

Write the ordered pair that represents vector YZ. Y(-2, 5), Z(1, 3) and write the magnitude

OpenStudy (zzr0ck3r):

you cant multiply vectors? is it dot or cross?

OpenStudy (anonymous):

i get confused

OpenStudy (zzr0ck3r):

well are you studing cross product or dot product?

OpenStudy (anonymous):

the ordered pair of YZ isnt a multiple it is when one is the origin so if you bring one point to the origin the other will change by the same amount

OpenStudy (zzr0ck3r):

ahh

OpenStudy (anonymous):

so to bring the x value of vector Y to zero you add two so you add two to vector Z's x component. same with the y component. to get Y's y component you subtract 5 so do the same to Z's. so the ordered pair is (3,-2)

OpenStudy (anonymous):

ok i understand that

OpenStudy (anonymous):

the magnitude is the magnitude of that ordered pair from the origin so \[\sqrt{3^{2}+(-2)^{2}}\]

OpenStudy (anonymous):

because that is the formula for distance and also magnitude of a vector

OpenStudy (anonymous):

ok so helpful thanks!!!!!

OpenStudy (anonymous):

no problem :) glad to help

OpenStudy (anonymous):

Find an ordered pair to represent vector a if vector b = <6, 3> and vector c = <-4, 8>.

OpenStudy (anonymous):

a=2b+c

OpenStudy (anonymous):

how would i do this? if you don't mind helping again

OpenStudy (anonymous):

there is a scalar on vector b and adding vectors is adding components. so for example b=<b1,b2> and c =<c1,c2> so 2b=<2b1,2b2> and 2b+c = <2b1+c1, 2b2+c2>

OpenStudy (anonymous):

in your vectors b =<6,3> c=-4,8> so 2b = <12,6> and 2b+c = <12+-4,6+8> = <8,14> = a